Dehner Corporation uses a job-order costing system with a single plantwide predetermined overhead rate based on direct labor-hours. The company based its predetermined overhead rate for the current year on the following data: Total direct labor-hours 94,000 Total fixed manufacturing overhead cost $404,200 Variable manufacturing overhead per direct labor-hour $ 4.00 Recently, Job P951 was completed with the following characteristics: Number of units in the job 50 Total direct labor-hours 100 Direct materials $ 660 Direct labor cost $9,400 The unit product cost for Job P951 is closest to: (Round your intermediate calculations to 2 decimal places.)
Answer #1

Solution: $217.80
Working: Estimated total manufacturing overhead cost = Estimated total fixed manufacturing OH cost + (Estimated variable OH cost per unit of the allocation base * Estimated total amount of the allocation base) = $404,200 + (94,000 * $4) = 780,200
Predetermined overhead rate = [Estimated total manufacturing OH cost / Estimated total amount of the allocation base] = 780,200 / 94000 = $8.3 per direct labor-hour
Overhead applied to a specific job = (Predetermined OH rate * Amount of the allocation base incurred by the job) = $8.3 per direct labor-hour * 100 = $830

Direct materials   660     
Direct labor cost   9400     
Manufacturing OH applied   830     
Total cost of Job P951   10890     
unit product cost for Job P951 = 10,890 / 50 units = $217.80